### v1

# Mechanical strength investigations of the APPLE-X undulator using Fiber Bragg Grating strain measurements

Abstract
The SPARCLAB facility at the INFN laboratory in Frascati is being upgraded to accommodate a new user facility as part of the SABINA (Source of Advanced Beam Imaging for Novel Applications) project. The SPARC laboratory was set up to investigate the feasibility of an ultra-brilliant photoinjector and to perform FEL experiments. Over the years, upgrades, additional beamlines, and experiments have been added. One of the latest upgrades, to be completed in 2024, will include improvements to the main infrastructure, laser and radiofrequency systems, and the implementation of a new electron beam line for SABINA. The beamline dedicated to SABINA will be equipped with three APPLE-X undulators acting as amplifiers to deliver IR/THz radiation with photon pulses in the ps range, with energy of tens of µJ, and with the possibility of choosing between linear, circular, or elliptical polarisation. The APPLE-X, designed and built by KYMA S.p.a., guarantees the possibility to vary the gap amplitude between the magnet’s arrays and their relative phase. The entire system, from the mechanics to the kinematic subsystems, has been designed from scratch, and a structural analysis has been carried out to evaluate the production. The undulators were delivered to Frascati in 2023 and, in collaboration with ENEA, a further investigation campaign was launched on the mechanical parts in the vicinity of the permanent magnets, using strain measurements based on optical methods. Fiber Bragg Grating sensors proved to be perfectly suitable to the requirements of these tests, mainly because of their immunity to electromagnetic noise. They consist of a phase grating inscribed in the core of a single-mode fiber, whose Bragg-diffracted light propagates back along the fiber. If bonded to the mechanical structure, they can be used as strain sensors. Indeed, any deformation of the monitored mechanical parts causes a deformation of the grating with a consequent change of its pitch and a related variation of the Bragg-diffracted wavelength. By following the variations in the scattered spectrum, it is possible to perform strain measurements. One µstrain resolution is within the reach of the current technology. Using multiple FBGs applied at selected locations on the undulator, several measurements were made by following the different possible kinematics, such as opening and closing the gap between the magnetic arrays or changing the phase, but also by studying the quiescent response as a function of the ambient temperature. The results of these tests show that there is a clear deformation of the structure related to the temperature changes and magnetic forces, but the magnitude of this deformation is well within the tolerances required for the functionality of the undulator since they are compatible or lower with respect to the one calculated with the finite elements methods. The tests, therefore, confirm the reliability of the mechanical structure.
1 SABINA
SABINA [1] is a project of the National Laboratory of Frascati (LNF), of the Italian Institute for Nuclear Physics (INFN), co-funded by the Lazio Region, whose goal is to create two user facilities by consolidating the existing SPARC_LAB [2] infrastructure. An irradiation facility with a laser target line for vacuum optical samples for aerospace and, the one of interest for this article, a THz/MIR radiation line for carrying out non-invasive studies in various fields of research. The SPARC_LAB LINAC provides an electron beam with high-brightness and tunable energies between and . For the THz line, the beam is directed into a dedicated dogleg. Here, 3 APPLE-X undulators operate in self-amplified spontaneous emission (SASE) mode creating photon pulses with frequency in the range or µ, with variable polarization, picosecond duration and energy of tens of µJ. The THz user facility will be promoted among scientific and industrial disciplines. The non-invasive technique that such radiation line offers is usable in several fields for studying complex phenomena and investigating inhomogeneities on a millimeter and sub-millimeter scale. It is expected to find collaborators among automotive, electronics, pharmaceutical or archeometry.
2 Tests
The undulators were manufactured by Kyma S.p.a. [3]. They are APPLE-X type with an innovative and compact design. A successful structural analysis was performed on the critical elements, and when they arrived at LNF, an additional campaign of measurements was conducted to test potential mechanical deformation at different magnetic field values. The studies of interest in this document are carried out in collaboration with ENEA [4], using Fiber Bragg Grating (FBG) instrumentation, which performs strain measurement. An overview of the experimental setup will be provided, featuring the combination of these two technologies for the first time, followed by a description of the tests done to collect data under different conditions, to have a clearer and less ambiguous perspective, as FBG sensors are sensitive to deformations caused by both temperature and mechanical stress.
FBG
The FBG [5] consist of a diffraction grating inscribed periodically in an optical fiber. The reflected or transmitted spectrum of light passing through the fiber is determined by the selected period of such gratings. As the grating changes, the characteristic wavelength of the sensors shifts, providing clear information about the strain that has occurred. The ENEA Fiber Optic Sensors Laboratory provided 1 cm long Broptics OS 1500 Optical Sensors, that have been readout by a Micro Optics FBG instrument.
APPLE-X
The undulators [6] are designed to have four magnetic arrays and a gap drive mechanism. This allows to change the gap (intensity of the magnetic field) by moving all the arrays simultaneously or the phase (the polarity of the radiation) by moving longitudinally two arrays arranged diagonally. The gap range is ; the allowed phase shift range is . The mechanical parts of interest in this analysis are the supports of the permanent magnet composed by three aluminum plates right below them and the single interface plate towards the girder. Figure 1(a) shows the details.
The sensors were installed on the arrays that are allowed to move also longitudinally, made the assumption that they would be the most stressed ones. 28 sensors (figure 1(b)) were glued with Araldite 2014, to investigate 3 different behaviors:
environment: 1 FBG monitors the behaviour with respect to the environment’s temperature. It is on a face of the metal plate outside the APPLE-X, close to the temperature sensors
gap sensors: 20 FBG placed on the surfaces facing the inside of the device between adjacent plates, either in vertical or horizontal position. In this case the strain measure gives indication of the changing in the separation place, overestimated to be at maximum
strain sensors: 7 FBG fully adhesive to the surface. The measurement in this case would be related to a mechanical deformation of the plate itself.
The measurements are made by following the peak of the spectrum with respect to the value at the beginning of the test: it is a comparative test. The value of the wavelength, is then converted into a strain parameter following the rule that corresponds to µ, defined as well as equal to µ.
Environmental
The response to temperature variation is tested with a static acquisition over 22 hours. The T was measured to be about . Two temperature’s sensors are used to give both a measure of the environment around the undulator (environment) and a measure of the metal plate, with the sensor installed adherent to the metal plate (plate). This test was helpful to notice, through figure 2, that the gap sensors are the one suffering a more significant change, while the strain sensors show changes compatibles to the results obtained with FBG00, sensitive only to the thermal expansion of the metal.
Gap Opening
The test to study the effects of the different intensity of the magnetic field is done registering the data while changing the gap of the undulator. Here is presented the test performed moving from the minimum gap amplitude up to with steps. Figure 3 reports only the most relevant behavior for the discussion. Both plots can be observed as splitted in two parts due to the different forces that are playing a role inside the undulator: before and after of the gap amplitude. At smaller gaps the horizontal sensors from opposite arrays show a relative strain of similar magnitude but opposite sign; after the behavior addresses probably a movement of the central plate due to the combined effect of magnetic, gravitational and mechanical forces, since that sensors around the same plate are either stretched or compressed. The vertical sensors show a maximum stress at about probably due only to the magnetic force, after which they do not move anymore.
Phase Shift
This last test is performed to investigate the last possible movimentation that the undulator can do. Figure 4 shows the most relevant steps of the tests performed at the minimum gap amplitude, i.e. at the higher magnetic force. The strain is slightly worse than the other situations presented, especially for horizontal sensors, but reproducible.
3 Conclusions
The described tests are performed with a well-known technology such as FBG sensors applied in a new way to investigate the performance of a new mechanical structure such the one of the undulator. The results were important to confirm that the high magnetic field does not affect the mechanical structure. They highlighted overall the extreme sensitivity of this measurement system to both temperature changes and mechanical movements and stress. In detail, a distinction can be done between the gap and the strain sensors. The greatest deformation occurs for the former, which is at most , among all the tests. For them, it is considered the length of the sensor not adherent to the metal by over-estimating the space between the metal plates to be . The results show a variation of the gap itself of . The strain sensors show a deformation lower of a factor 2 to 5. Here, the deformation occurs over the full length of the sensors and can be estimated to be at worst of . The deformations are compatible or lower than the ones calculated by the finite elements method [6] and, therefore, confirm the reliability of the undulator mechanics.
Acknowledgments
This work is supported by NextGeneration EU- Italian National Recovery and Resilience Plan, Mission 4 - Component 2 - Investment 3.1. - Project name: Rome Technopole, CUP: I93C21000150006. SABINA is a project co-funded by Regione Lazio with the “Research infrastructures” public call within PORFESR 2014-2020.